# Differentially Private Zeroth-Order Methods for Scalable Large Language Model Fine-tuning

### v4

@@ 3.3. Zeroth-Order Optimization
Definition 3.8 (Zeroth-order gradient estimation (Spall, 1992)).
@@
(8) $\nabla f_{\beta}(\boldsymbol{\theta},\xi)=\frac{1}{2\beta}\left(f(\boldsymbol{\theta}+\beta\mathbf{v},\xi)-f(\boldsymbol{\theta}-\beta\mathbf{v},\xi)\right)\mathbf{v},$
where $\mathbf{v}\in\mathcal{R}^{d}$ with $\mathbf{v}\sim N(0,\textbf{I}_{d})$ and $\beta$ is the ZO scale parameter.
Gaussian Smoothing: Gaussian smoothing is a well-known technique that converts a possibly non-smooth function to a smooth approximation (Nesterov and Spokoiny, 2017). Given a differentiable function $f:\mathcal{R}^{d}\rightarrow\mathcal{R}$ and $\beta\geq 0$, we define the Gaussian smoothing of $f$ as $f_{\beta}(\boldsymbol{\theta})=\mathbb{E}_{\mathbf{v}\sim N(0,\textbf{I}_{d})}\left[f(\boldsymbol{\theta}+\beta\mathbf{v})\right]$ where $N(0,\textbf{I}_{d})$ is a standard Gaussian distribution. Gaussian smoothing has the following properties.
Lemma 3.9.
@@
(9) $\nabla f_{\beta}(\boldsymbol{\theta})=\mathbb{E}_{\mathbf{v}\sim N(0,\textbf{I}_{d})}\left[\frac{1}{2\beta}\left(f(\boldsymbol{\theta}+\beta\mathbf{v})-f(\boldsymbol{\theta}-\beta\mathbf{v})\right)\mathbf{v}\right].$
Weakly-convex function is a class of “nice-behaviored” non-convex objectives that are close to a convex function which is well studied as non-convexity assumption in (Yuan et al., 2019) (Bassily et al., 2021).
Definition 3.10 (Weakly-convex).
@@
(10) $f(\boldsymbol{\theta}^{\prime})\geq f(\boldsymbol{\theta})+\langle\nabla f(\boldsymbol{\theta}),\boldsymbol{\theta}^{\prime}-\boldsymbol{\theta}\rangle-\frac{\rho}{2}\left\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\right\|^{2}$
@@
The function $f_{\beta}(\boldsymbol{\theta},x)$ is $L$-Lipschitz and $\frac{\sqrt{d}L}{\beta}$-smooth for every $x$. There exist $\gamma$ that for every $x\in\mathcal{D}$, we have
(11) $\mathbb{E}\left[\left\|\nabla F_{\beta}(\boldsymbol{\theta})-\nabla f_{\beta}(\boldsymbol{\theta},x)\right\|^{2}\right]\leq\gamma^{2}.$
It is notable that many papers have proposed and analyzed deterministic/stochastic optimization algorithms under the PL condition (Karimi et al., 2016) (Bassily et al., 2018).
@@
(12) $\left\|\boldsymbol{\theta}-\boldsymbol{\theta}^{*}\right\|^{2}\leq\frac{1}{2\mu}\left(f_{\beta}(\boldsymbol{\theta},\mathcal{D})-f_{\beta}(\boldsymbol{\theta}^{*},\mathcal{D})\right).$
4. Methodology
In Section 4.1, we introduce differential private zeroth-order fine-tuning and describe our novel DP noise injection method for zeroth-order. In Section 5.7, we propose zeroth-order gradient estimation with dynamic ZO scale parameter. In Section 4.3, we propose DP-ZOSO, the first stagewise DP-ZO fine-tuning method to our knowledge, and provide theoretical analysis for both privacy and convergence. In Section 4.4, we propose DP-ZOPO, the first stagewise DP-ZO fine-tuning method with data-free pruning to guide the model to update on important directions, whose schematic diagram is shown in Figure LABEL:fig_pruning.
4.1. Differential Private Zeroth-Order Fine-Tuning
@@
Compared to adding Gaussian to all dimensions of the gradient in DP-SGD, we propose a novel method of noise injection by adding noise $N(0,\sigma^{2}C^{2})$ to $\frac{1}{2\beta}\left(f(\boldsymbol{\theta}+\beta\mathbf{v})-f(\boldsymbol{\theta}-\beta\mathbf{v})\right)$, achieving $(\epsilon,\delta)$-DP guarantee. By standard post processing, the approximate gradient $\nabla f_{\beta}(\boldsymbol{\theta})$ is also $(\epsilon,\delta)$-DP since the direction $\mathbf{v}$ is independent of private data.
4.2. Zeroth-Order Gradient Estimation with Dynamic ZO scale parameter
The zeroth-order method suffers from utility loss owing to the bias between the actual gradient and the zeroth-order gradient. Zeroth-order gradient is estimated by calculating the difference of the function on two points $\boldsymbol{\theta}+\beta\mathbf{v}$ and $\boldsymbol{\theta}-\beta\mathbf{v}$ where $\beta$ is the ZO scale parameter. The choice of ZO scale parameter $\beta$ has only been seen as a hyperparameter in (Malladi et al., 2023) (Ghadimi and Lan, 2013). Shi et al. (Shi et al., 2022) pointed out that the ZO scale parameter should be as small as possible but can be set too small in practice since the accuracy of the computing system is limited. However, when the model is close to convergence, the difference of the loss function on two points can be quite small such that the two-point estimated gradient can be quite large when the ZO scale parameter is small, contrary to the nature of the model approaching convergence. Thus, we propose dynamic zeroth-order gradient estimation with an increasing ZO scale parameter, reducing the bias of the zeroth-order gradient. More results are detailed in Table 5.
Furthermore, we find that zeroth-order gradients with bigger ZO scale parameters have. The lower the possibility of zeroth-order gradient getting clipped (detailed in Table 1).
Table 1. Clipping bias with different ZO scale parameters. $\mathbb{P}(\mathrm{Clip})$ denotes the possibility of getting clipped.
@@ 4.2. Zeroth-Order Gradient Estimation with Dynamic ZO scale parameter
4.3. DP-ZO Stagewise Fine-Tuning
First, let us propose algorithm Differentially Private Zeroth-Order Stagewise Optimizer (DP-ZOSO) in Algorithm 1. At the $s$-th stage, a regularized function $\phi_{s}(\boldsymbol{\theta})$ is constructed that consists of the original objective $f_{\beta_{s}}(\boldsymbol{\theta})$ and a quadratic regularizer $\frac{1}{2\lambda}\left\|\boldsymbol{\theta}-\boldsymbol{\theta}^{s-1}\right\|_{2}^{2}$. The reference point $\boldsymbol{\theta}^{s-1}$ is a returned solution from the previous stage, which is also used for an initial solution for the current stage. Adding the strongly convex regularizer at each stage helps convert the weakly-convex loss function $f_{\beta_{s}}(\boldsymbol{\theta})$ to convex or strongly-convex $\phi_{s}(\boldsymbol{\theta})$. For each regularized problem, the SGD with a constant stepsize and ZO scale parameter is employed for several iterations with an appropriate returned solution. The procedure is described in Algorithm 2.
input : Initial point $\boldsymbol{\theta}^{0}\in\mathcal{R}^{d}$, initial stepsize $\eta_{0}\geq 0$, initial ZO scale parameter $\beta_{0}$, regularization parameter $\lambda$, initial iteration number $T_{0}$, number of epoch $S$, dataset $\mathcal{D}$
@@ 4.3. DP-ZO Stagewise Fine-Tuning
2       $\beta_{s}=k\cdot\beta_{s-1},\ T_{s}=2\cdot T_{s-1},\ \eta_{s}=\eta_{s-1}/2,\ \lambda=\lambda_{s-1}/2$;
@@
Algorithm 2 is a Differentially Private Zeroth-Order Optimizer (DP-ZOO) that updates parameters in specific distribution $\mathbb{V}$. In the step $t$ of Algorithm 2, a random set of $\xi_{t}$ is sampled from dataset $\mathcal{D}$. Random $P$ vectors $\left\{\mathbf{v}_{t}^{p}\right\}_{p=1}^{P}\in\mathbb{V}$ are sampled from distribution $\mathbb{V}$. For vector $\mathbf{v}_{t}^{p}$, we estimate the gradient step $\mathrm{gra}_{f,t}^{i,p}$ of original objective $f(\boldsymbol{\theta}_{t-1})$ on every data point $x_{t}^{i}$ in line 7 and then clip it to $\hat{\mathrm{gra}}_{f,t}^{i,p}$ in L2-norm. The gradient step of the quadratic regularizer is added to the clipped gradient step as $\mathrm{gra}_{t}^{i,p}$ in line 9. Gaussian noise is added to the sum of $\mathrm{gra}_{t}^{i,p}$ and then average. Timing the direction $\mathbf{v}_{t}^{p}$ is the gradient $g_{p}(\boldsymbol{\theta}_{t-1})$ on direction $\mathbf{v}_{t}^{p}$. The gradient in iteration $t$ is $g(\boldsymbol{\theta}_{t-1})$ by taking average over $P$ directions. Parameters are then updated by $g(\boldsymbol{\theta}_{t-1})$ with learning rate $\eta_{s}$.
@@ 4       Sample $P$ random vectors $\footnotesize\left\{\mathbf{v}_{t}^{p}\right\}_{p=1}^{P}\in\mathbb{V}$;
7                   $\footnotesize\mathrm{gra}_{f,t}^{i,p}=\frac{1}{2\beta_{s}}\left(f(\boldsymbol{\theta}_{t-1}+\beta_{s}\mathbf{v}_{t}^{p},x_{t}^{i})-f(\boldsymbol{\theta}_{t-1}-\beta_{s}\mathbf{v}_{t}^{p},x_{t}^{i})\right)$ ;
@@
9                   $\mathrm{gra}_{t}^{i,p}=\hat{\mathrm{gra}}_{f,t}^{i,p}+\frac{1}{\lambda}\left\|\boldsymbol{\theta}_{t-1}-\boldsymbol{\theta}_{0}\right\|$;
10
@@
12            $g_{p}(\boldsymbol{\theta}_{t-1})=\frac{1}{m}\left(\sum_{i=1}^{m}\mathrm{gra}_{t}^{i,p}+N(0,\sigma^{2}C^{2})\right)\cdot\mathbf{v}_{t}^{p}$;
13
@@ 16       $\boldsymbol{\theta}_{t}=\boldsymbol{\theta}_{t-1}-\eta_{s}\cdot g(\boldsymbol{\theta}_{t-1})$;
Theorem 4.1 (Privacy analysis of DP-ZOSO).
@@
(13) $\sigma^{2}\geq\frac{c_{2}^{2}P^{2}m^{2}T\log(P/\delta)}{\epsilon^{2}n^{2}}.$
Proof.
In the algorithm, the sensitivity of the sum of ZO estimated $\sum_{i=1}^{m}\hat{\mathrm{gra}}_{f,t}^{i,p}$ is clipped to $C$. The zeroth-order gradient on data point $x_{t}^{i}$ is estimated in $P$ random directions such that the privacy budget is divided equally into $P$. Vectors are randomly sampled from standard Gaussian distribution, thus by post-processing, the privacy of the zeroth-order gradient is guaranteed. ∎
Lemma 4.2.
@@
(14) $\begin{split}\mathbb{E}[||\nabla F_{\beta}(\boldsymbol{\theta}_{t})-&\nabla\hat{f}_{\beta}(\boldsymbol{\theta}_{t},\xi_{t+1})||^{2}]\\ \leq&\frac{dc_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{8dC^{2}}{ePm}+\frac{64d\beta^{2}L^{4}}{Pm}+\frac{\gamma^{2}}{Pm},\end{split}$
where the first two terms are due to DP noise and clipping and the last two terms are due to zeroth-order error and data sampling error.
Lemma 4.3.
@@
(15) $\begin{split}\mathbb{E}&[\phi_{s}(\boldsymbol{\theta}_{T_{s}})-\phi_{s}(\boldsymbol{\theta})]\leq\left(\frac{1}{2\eta_{s}T_{s}}+\frac{1}{2T_{s}\lambda}\right)\left\|\boldsymbol{\theta}_{0}-\boldsymbol{\theta}\right\|^{2}\\ &+\eta_{s}\left(\frac{dc_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{64d\beta_{s}^{2}L^{4}}{Pm}+\frac{\gamma^{2}}{Pm}+\frac{8dC^{2}}{ePm}\right).\\ \end{split}$
Theorem 4.4.
@@
(16) $\phi_{s}(\boldsymbol{\theta}^{S})-\phi_{s}(\boldsymbol{\theta}^{*})\leq\alpha.$
@@
(17) $\mathcal{O}\left(\left(\gamma^{2}+dC^{2}+d\beta_{S}^{2}L^{4}+\frac{dC^{2}P^{2}m\log(P/\delta)}{\epsilon^{2}n^{2}}\right)\cdot\frac{1}{\mu\alpha}\right).$
4.4. DP-ZO Stagewise Fine-Tuning with pruning
The success in parameter-efficient fine-tuning motivates us that fine-tune a subset of parameters LLMs can significantly improve the performance on downstream tasks. Pruning helps the model to update on more important directions, especially in the setting of differentially private zeroth-order fine-tuning. Compared with first-order optimizer, zeroth-order optimizer in conjunction with pruning yields greater enhancement in private settings (more details in Section 5.5).

### v5

@@ III-C Zeroth-Order Optimization
Definition 5 (Zeroth-order gradient estimation [35]).
@@
$\nabla f_{\beta}(\boldsymbol{\theta},\xi)=\frac{1}{2\beta}\left(f(\boldsymbol{\theta}+\beta\mathbf{v},\xi)-f(\boldsymbol{\theta}-\beta\mathbf{v},\xi)\right)\mathbf{v},$ (8)
where $\mathbf{v}\sim N(0,\textbf{I}_{d})$ and $\beta$ is the ZO scale parameter.
Gaussian Smoothing: Gaussian smoothing is a well-known technique that converts a possibly non-smooth function to a smooth approximation [36]. Given a differentiable function $f:\mathcal{R}^{d}\rightarrow\mathcal{R}$ and $\beta\geq 0$, we define the Gaussian smoothing of $f$ as $f_{\beta}(\boldsymbol{\theta})=\mathbb{E}_{\mathbf{v}\sim N(0,\textbf{I}_{d})}\left[f(\boldsymbol{\theta}+\beta\mathbf{v})\right]$ where $N(0,\textbf{I}_{d})$ is a standard Gaussian distribution. Gaussian smoothing has the following properties.
Lemma 3.
@@
$\nabla f_{\beta}(\boldsymbol{\theta})=\mathbb{E}_{\mathbf{v}\sim N(0,\textbf{I}_{d})}\left[\frac{1}{2\beta}\left(f(\boldsymbol{\theta}+\beta\mathbf{v})-f(\boldsymbol{\theta}-\beta\mathbf{v})\right)\mathbf{v}\right].$ (9)
Weakly-convex function is a class of “nice-behaviored” non-convex objectives that are close to a convex function which is well studied as non-convexity assumption in [37] [33].
Definition 6 (Weakly-convex).
@@
$f(\boldsymbol{\theta}^{\prime})\geq f(\boldsymbol{\theta})+\langle\nabla f(\boldsymbol{\theta}),\boldsymbol{\theta}^{\prime}-\boldsymbol{\theta}\rangle-\frac{\rho}{2}\left\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\right\|^{2}$ (10)
@@
The function $f_{\beta}(\boldsymbol{\theta},x)$ is $L$-Lipschitz and $\frac{\sqrt{d}L}{\beta}$-smooth. There exist $\gamma$ that for every $x\in\mathcal{D}$, we have
$\mathbb{E}\left[\left\|\nabla F_{\beta}(\boldsymbol{\theta})-\nabla f_{\beta}(\boldsymbol{\theta},x)\right\|^{2}\right]\leq\gamma^{2}.$ (11)
It is notable that many papers have proposed and analyzed deterministic/stochastic optimization algorithms under the PL condition [38] [39].
@@
$\left\|\boldsymbol{\theta}-\boldsymbol{\theta}^{*}\right\|^{2}\leq\frac{1}{2\mu}\left(f_{\beta}(\boldsymbol{\theta},\mathcal{D})-f_{\beta}(\boldsymbol{\theta}^{*},\mathcal{D})\right).$ (12)
IV Methodology
In Section IV-A, we introduce differential private zeroth-order fine-tuning and describe our novel DP noise injection method for zeroth-order. In Section IV-B, we propose zeroth-order gradient estimation with dynamic ZO scale parameter. In Section IV-C, we propose DP-ZOSO, the first stagewise DP-ZO fine-tuning method to our knowledge, and provide theoretical analysis for both privacy and convergence. In Section IV-D, we propose DP-ZOPO, the first stagewise DP-ZO fine-tuning method with data-free pruning to guide the model to update on important directions, whose schematic diagram is shown in Figure LABEL:fig_pruning.
IV-A Differential Private Zeroth-Order Fine-Tuning
@@
Compared to adding Gaussian to all dimensions of the gradient in DP-SGD, we propose a novel method of noise injection by adding noise $N(0,\sigma^{2}C^{2})$ to $\frac{1}{2\beta}\left(f(\boldsymbol{\theta}+\beta\mathbf{v})-f(\boldsymbol{\theta}-\beta\mathbf{v})\right)$, achieving $(\epsilon,\delta)$-DP guarantee. By standard post processing, the approximate gradient $\nabla f_{\beta}(\boldsymbol{\theta})$ is also $(\epsilon,\delta)$-DP since the direction $\mathbf{v}$ is independent of private data. We also derive experimental results that with straightforward DP noise injection, the model fails to converge and exhibits output patterns almost indistinguishable from random noise.
IV-B Zeroth-Order Gradient Estimation with Dynamic ZO scale parameter
The zeroth-order method suffers from utility loss owing to the bias between the actual gradient and the zeroth-order gradient. Zeroth-order gradient is estimated by calculating the difference of the function on two points $\boldsymbol{\theta}+\beta\mathbf{v}$ and $\boldsymbol{\theta}-\beta\mathbf{v}$ where $\beta$ is the ZO scale parameter. The choice of ZO scale parameter $\beta$ has only been seen as a hyperparameter in [20] [40]. Shi et al. [41] pointed out that the ZO scale parameter should be as small as possible but can be set too small in practice since the accuracy of the computing system is limited. However, when the model is close to convergence, the difference of the loss function on two points can be quite small such that the two-point estimated gradient appears disproportionately large when the ZO scale parameter is small. Thus, we propose dynamic zeroth-order gradient estimation with an increasing ZO scale parameter, reducing the bias of the zeroth-order gradient and offering a stabilization effect.
Furthermore, we find that zeroth-order gradients with bigger ZO scale parameters have the lower possibility of zeroth-order gradient getting clipped (detailed in Table I). More results of training loss with fix and dynamic ZO scale parameter are detailed in Table IV.
TABLE I: Clipping bias of RoBERTa-large fine-tuned on 1000 examples of SNLI with different ZO scale parameters. $\mathbb{P}(\mathrm{Clip})$ denotes the possibility of getting clipped.
@@ IV-B Zeroth-Order Gradient Estimation with Dynamic ZO scale parameter
IV-C DP-ZO Stagewise Fine-Tuning
First, let us propose algorithm Differentially Private Zeroth-Order Stagewise Optimizer (DP-ZOSO) in Algorithm 1. We present the algorithm workflow in Figure LABEL:fig_algorithm. The training process is divided into $S$ stages. At the $s$-th stage, the regularization term $\frac{1}{2\lambda}\left\|\boldsymbol{\theta}-\boldsymbol{\theta}^{s-1}\right\|_{2}^{2}$ plays a crucial role in converting the weakly-convex loss function $f_{\beta_{s}}(\boldsymbol{\theta})$ to convex or strongly-convex $\phi_{s}(\boldsymbol{\theta})$, which divides the more difficult non-convex objective into easier and more tractable ones for each stage. The regularized function $\phi_{s}(\boldsymbol{\theta})$ is updated with the reference point $\boldsymbol{\theta}^{s-1}$ which is the returned solution from the previous stage. In earlier stages, the gradient approximation error is controlled to be smaller with smaller ZO scale parameter, and together with a larger learning rate, it allows the LLM fine-tuning to approach the optimum more quickly. In the later stages, we need to deliberately increase the gradient approximation error, together with a decreased learning rate, offering a stabilization effect on the fine-tuning trajectory.
input : Initial point $\boldsymbol{\theta}^{0}\in\mathcal{R}^{d}$, initial stepsize $\eta_{0}\geq 0$, initial ZO scale parameter $\beta_{0}$, regularization parameter $\lambda$, initial iteration number $T_{0}$, number of stage $S$, dataset $\mathcal{D}$
@@ IV-C DP-ZO Stagewise Fine-Tuning
2       $\beta_{s}=k\cdot\beta_{s-1},\ T_{s}=2\cdot T_{s-1},\ \eta_{s}=\eta_{s-1}/2$;
@@
Algorithm 2 is a Differentially Private Zeroth-Order Optimizer (DP-ZOO) that updates parameters in specific distribution $\mathbb{V}$. In the step $t$ of Algorithm 2, $\xi_{t}$ is uniformly sampled without replacement from dataset $\mathcal{D}$. Random $P$ vectors $\left\{\mathbf{v}_{t}^{p}\right\}_{p=1}^{P}\in\mathbb{V}$ are sampled from distribution $\mathbb{V}$. For vector $\mathbf{v}_{t}^{p}$, we estimate the gradient step $\mathrm{gra}_{f,t}^{i,p}$ of original objective $f(\boldsymbol{\theta}_{t-1})$ on every data point $x_{t}^{i}$ in line 7 and then clip it to $\hat{\mathrm{gra}}_{f,t}^{i,p}$ in L2-norm. Gaussian noise is added to the sum of $\hat{\mathrm{gra}}_{f,t}^{i,p}$ and then average. Timing the direction $\mathbf{v}_{t}^{p}$ is the DP zeroth-order gradient $g_{p}(\boldsymbol{\theta}_{t-1})$ on direction $\mathbf{v}_{t}^{p}$. The gradient of the quadratic regularizer is added to the noisy clipped gradient in line 11 (no need to add DP noise). The gradient in iteration $t$ is $g(\boldsymbol{\theta}_{t-1})$ by taking average over $P$ directions. Parameters are then updated by $g(\boldsymbol{\theta}_{t-1})$ with learning rate $\eta_{s}$. We choose to add DP on the clipped ZO gradients of the original loss $\hat{\mathrm{gra}}_{f,t}^{i,p}$, because the varying gradient of the regularizer $\frac{1}{\lambda}(\boldsymbol{\theta}_{t-1}-\boldsymbol{\theta}_{0})$ does not access private data which will not cause privacy leakage and will complicate the clipping threshold setting.
@@ 4       Sample $P$ random vectors $\footnotesize\left\{\mathbf{v}_{t}^{p}\right\}_{p=1}^{P}\in\mathbb{V}$;
7                   $\scriptsize\mathrm{gra}_{f,t}^{i,p}=\frac{1}{2\beta_{s}}\left(f(\boldsymbol{\theta}_{t-1}+\beta_{s}\mathbf{v}_{t}^{p},x_{t}^{i})-f(\boldsymbol{\theta}_{t-1}-\beta_{s}\mathbf{v}_{t}^{p},x_{t}^{i})\right)$ ;
@@
9
@@
11            $\small g_{p}(\boldsymbol{\theta}_{t-1})=\frac{1}{m}\left(\sum_{i=1}^{m}\hat{\mathrm{gra}}_{f,t}^{i,p}+N(0,\sigma^{2}C^{2})\right)\cdot\mathbf{v}_{t}^{p}$;
12             $g_{p}(\boldsymbol{\theta}_{t-1})=g_{p}(\boldsymbol{\theta}_{t-1})+\frac{1}{\lambda}\left(\boldsymbol{\theta}_{t-1}-\boldsymbol{\theta}_{0}\right)$
@@ 15       $\boldsymbol{\theta}_{t}=\boldsymbol{\theta}_{t-1}-\eta_{s}\cdot g(\boldsymbol{\theta}_{t-1})$;
Theorem 1 (Privacy analysis of DP-ZOSO).
@@
$\sigma^{2}\geq\frac{c_{2}^{2}P^{2}m^{2}T\log(P/\delta)}{\epsilon^{2}n^{2}}.$ (13)
Proof of Theorem 13.
In DP-ZOSO, the sensitivity of every ZO estimated $\hat{\mathrm{gra}}_{f,t}^{i,p}(i\in[1,m])$ is clipped to $C$. The zeroth-order gradient on data point $x_{t}^{i}$ is estimated in $P$ random directions such that the privacy budget is divided equally into $P$. Since vectors are randomly sampled from standard Gaussian distribution, thus by post-processing, it suffices to show that the privacy of the term $\frac{1}{m}\sum_{i=1}^{m}\hat{\mathrm{gra}}_{f,t}^{i,p}$ is guaranteed. By applying Lemma 1 and Lemma 2, DP-ZOSO satisfies $(\epsilon,\delta)$-DP if $\sigma^{2}\geq\frac{c_{2}^{2}P^{2}m^{2}T\log(P/\delta)}{\epsilon^{2}n^{2}}$. ∎
Before we analyze the convergence of DP-ZOSO, we will first present the approximation bound for noisy ZO gradients in the following Lemma 4.
Lemma 4.
@@
$\begin{split}\mathbb{E}[||\nabla&F_{\beta}(\boldsymbol{\theta}_{t})-\nabla\hat{f}_{\beta}(\boldsymbol{\theta}_{t},\xi_{t+1})||^{2}]\\ \leq&\frac{dc_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{8dC^{2}}{ePm}+\frac{64d\beta^{2}L^{4}}{Pm}+\frac{\gamma^{2}}{Pm},\end{split}$ (14)
where the first two terms are due to DP noise and clipping bias and the last two are due to zeroth-order estimation and data sampling errors.
Lemma 5 presents the utility of DP-ZOO which is used in Theorem 2 to derive the total ZO oracle complexity.
Lemma 5.
@@
$\begin{split}\mathbb{E}&[\phi_{s}(\boldsymbol{\theta}_{T_{s}})-\phi_{s}(\boldsymbol{\theta})]\leq\left(\frac{1}{2\eta_{s}T_{s}}+\frac{1}{2T_{s}\lambda}\right)\left\|\boldsymbol{\theta}_{0}-\boldsymbol{\theta}\right\|^{2}\\ &+\eta_{s}\left(\frac{dc_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{64d\beta_{s}^{2}L^{4}}{Pm}+\frac{\gamma^{2}}{Pm}+\frac{8dC^{2}}{ePm}\right).\\ \end{split}$ (15)
Theorem 2.
@@
$\phi_{s}(\boldsymbol{\theta}^{S})-\phi_{s}(\boldsymbol{\theta}^{*})\leq\alpha.$ (16)
@@
$\mathcal{O}\left(\left(\gamma^{2}+dC^{2}+d\beta_{S}^{2}L^{4}+\frac{dC^{2}P^{2}m\log(P/\delta)}{\epsilon^{2}n^{2}}\right)\cdot\frac{1}{\mu\alpha}\right).$ (17)
IV-D DP-ZO Stagewise Fine-Tuning with pruning
The success in parameter-efficient fine-tuning motivates us that fine-tuning a subset of parameters LLMs can significantly improve the performance on downstream tasks. Pruning helps the model to update on more important directions, especially in the setting of differentially private zeroth-order fine-tuning. Compared with first-order optimizer, zeroth-order optimizer in conjunction with pruning yields greater enhancement in private settings (more details in Section V-E).